OpenStudy (anonymous):

prove that 2+cos^2x-3cos^4x=sin^2x(2+3cos^2x)

OpenStudy (jdoe0001):

\(\bf recall\implies sin^2(\theta)+cos^2(\theta)=1\implies {\color{blue}{ sin^2(\theta)}}=1-cos^2(\theta)\\ \quad \\ \quad \\ 2+cos^2(x)-3cos^4(x)={\color{blue}{ sin^2(x)}}[2+3cos^2(x)]\)

OpenStudy (anonymous):

Thank you!!